OpenStudy (alex344344):

What are the solutions to the equation (2x – 5)(3x – 1) = 0? x = mc003-1.jpg or x = mc003-2.jpg x = mc003-3.jpg or x = 3 x = mc003-4.jpg or x = mc003-5.jpg x = 5 or x = 1

OpenStudy (fwizbang):

You are multiplying two numbers, 2x-5 and 3x-1, ang getting zero. The only way that can happen is if one of the two numbers is zero....

OpenStudy (queenie):

2x - 5 = 0 2x = 5 x = 5/2 3x - 1 = 0 3x = 1 x = 1/3

OpenStudy (alex344344):

thanksss